OpenStudy (foreverandeverandeverrrr021014):

Let f(x) = x2 − 8x + 5. Find f(−1). -3 -4 14 15. its def not -3 please help

OpenStudy (anonymous):

plug in -1 into x

OpenStudy (anonymous):

-2+8+5

OpenStudy (janu16):

yea do that and solve it and you should end up with 14

OpenStudy (dinamix):

f(-1)=1+8+5=14 @foreverandeverandeverrrr021014

OpenStudy (foreverandeverandeverrrr021014):

thank you !

OpenStudy (dinamix):

i'm sure 200%

OpenStudy (anonymous):

the options are incorrect

OpenStudy (janu16):

the options are correct

OpenStudy (anonymous):

sorry 12

OpenStudy (mathmale):

Let f(x) = x2 − 8x + 5. Find f(−1). Please type "the square of x" as x^2. Secondly, please double-check on something else: Are you sure you're supposed to find f(-1), or is it\[f ^{-1}(x)?\]